OpenStudy (anonymous):

Coach Johnson is 10 years older than Tim. When Tim started football 3 years ago, Coach Johnson was twice his age. How old is Tim now? explain

OpenStudy (anonymous):

Ask Tim for his age and he will tell you. :) I hate these kinds of problems.

OpenStudy (mathmagician):

you have a system of equations: x=10+y x-3=2(y-3) =>(10+y)-3=2y-6 =>-y=-13=>y=13 and x =23. Now Toms age is 13 years.

OpenStudy (anonymous):

J be coach Johnson's age now. T be Tim's age now. We know: J = T + 10 We also know: J - 3 = 2(T-3) Substitute and solve for T as follows: T + 10 - 3 = 2(T-3) T + 7 = 2T - 6 T = 13 Years